Father's age is three years more than three times the son's age. After three years, father's age will be ten years more than twice the son's age. What is the father's present age.
Answer by rapaljer(4671)

Let x = son's present age
3x+3 = father's present age

x+3 = son's age in 3 years
3x+3 + 3 = father's age in 3 years

"Father's age will be ten more than twice the son's age"
3x+3 + 3 = 2 (x+3) + 10
3x + 6 = 2x + 6 + 10
x = 10 = son's age now
3x+3 = 33 = father's age now

Check: In 3 years, the son will be 13 and the father will be 36, which is indeed 10 more than twice the son's age.

R^2 from SCC